# Recursive subhomogeneity of orbit-breaking subalgebras of $\mathrm{C}^{*}$-algebras associated to minimal homeomorphisms twisted by line bundles

### v1

@@ 7. Embedding $\mathcal{O}(\mathcal{E}_{Y})$ into a direct sum of homogeneous algebras
Let $\varsigma=(\varsigma_{1},\ldots,\varsigma_{K})\in\Gamma(\mathscr{M}_{1})\oplus\cdots\oplus\Gamma(\mathscr{M}_{K})$. For an integer $10\leq m\leq r_{K}-$ 1, we say that $\varsigma_{k}\in\Gamma(\mathscr{M}_{k})$ has an $m$th subdiagonal matrix representation if at every $x\in\overline{Y_{k}}$, the matrix $(H_{U}^{(k)})^{-1}(x,\varsigma_{k}(x))$ is $m$th subdiagonal for some (hence for all) $U\in\mathcal{U}$ with $x\in U$. Here $(H_{U}^{(k)})^{-1}(x,\varsigma_{k}(x))$ is regarded as the zero matrix if $m\geq r_{k}$.
Assume that each $\varsigma_{k}\in\Gamma(\mathscr{M}_{k})$ has an $m$th lower subdiagonal matrix representation. We know from Remark 6.2 that for $x\in\overline{Y_{k}}$, the map
@@
From the definition of $\pi_{k}$, we know that if $g\in C(X)$, then $\pi_{k}(g)(x)$ is represented as a diagonal matrix for all $x\in\overline{Y_{k}}$. Also if $\xi\in\mathcal{E}_{Y}^{\otimes m}\subset\mathcal{O}(\mathcal{E}_{Y})$ for some $11\leq m\leq r_{k}-$ 1, then $\pi_{k}(\xi)(x)$ is represented as an $m$th lower subdiagonal matrix since it maps $\mathscr{V}_{x}^{(i)}$ into $\mathscr{V}_{x}^{(i+m)}$ by (21). Also if $\varsigma=(\varsigma_{1},\ldots,\varsigma_{K})=\pi(\xi)$ for $\xi\in\mathcal{E}^{\otimes m}$, then we have $\varsigma_{k}^{(i)}(x)=\psi(\xi)(\alpha^{i}(x))$ by (21) and (14).
@@
We assume $r_{k}>1$. Since the $C^{*}$-algebra $\Gamma(\mathscr{M}_{k})$ is generated by the sections $\varsigma\in\Gamma(\mathscr{M}_{k})$ which is an $m$th lower subdiagonal matrix in its matrix representation for $0\leq m<r_{k}$, it is enough to show that $\varsigma\in\pi_{k}(\mathcal{O}(\mathcal{E}_{Y}))$ for any such section $\varsigma$. Note that if $0\leq i\neq j<r_{k}$, then $\alpha^{i}(Y_{k})\cap\alpha^{j}(Y_{k})=\emptyset$ by definition of $r_{k}$.
Let $\varsigma\in\Gamma(\mathscr{M}_{k})$ have $m$th lower subdiagonal matrix representation. Then we know from the discussion given before this lemma that for $i=0,\dots,r_{k}-m-1$, there exists $\varsigma^{(i)}(x)\in\mathscr{V}_{\alpha^{i}(x)}^{(m)}$ such that the map
is given by $v\mapsto\varsigma^{(i)}(x)\otimes v$ for $v\in\mathscr{V}^{(i)}_{x}$.
@@
$\xi|_{\alpha^{i}(Y_{k})}(\alpha^{i}(x)):=\varsigma^{(i)}(x)$
for $x\in Y_{k}$ and $i=0,\dots,r_{k}-m-1$. This $\xi$ is well-defined since the sets $\alpha^{i}(Y_{k})$ are mutually disjoint. It is easy to see that
@@
and then by the Tietze extension theorem for a vector bundle (see [5, Section 1.4]), $\xi$ extends to a section $\widetilde{\xi}\in\Gamma(\mathscr{V}^{(m)})$ over the whole space $X$ such that
@@
$=$ $\ (0,\ldots,0,\varsigma^{(0)}(x)\otimes a_{0},\varsigma^{(1)}(x)\otimes a_{1},\ldots,\varsigma^{(r_{k}-m-1)}(x)\otimes a_{r_{k}-m-1})$
$=$ $\ \varsigma(x)(a_{0},a_{1},\ldots,a_{r_{k}-1}),$
hence $\pi_{k}(\psi^{-1}(\widetilde{\xi}))=\varsigma$ follows. ∎
@@ 8. A recursive subhomogeneous decomposition of $\mathcal{O}(\mathcal{E}_{Y})$
In this section we explicitly describe the image of $\mathcal{O}(\mathcal{E}_{Y})$ under the injective -homomorphism $\pi:\mathcal{O}(\mathcal{E}_{Y})\to\Gamma(\mathscr{M}_{1})\oplus\dots\oplus\Gamma(\mathscr{M}_{K})$. In particular, we construct a recursive subhomogeneous decomposition for the range of $\pi$ which yields that the orbit-breaking algebra $\mathcal{O}(\mathcal{E}_{Y})$ is a recursive subhomogeneous algebra. As in the case of $\mathrm{C}^{*}$-algebras of minimal homeomorphisms, we investigate the behaviour of boundary points. However, the presence of a twist (that is, a non-trivial line bundle) leads to further difficulties since we need to consider the local trivialization of the underlying endomorphism bundles when we are tracking the travel of boundary points through the Rokhlin towers.
@@
for $x\in\overline{Y_{k}}$, $2\leq k\leq K$, Note that the vector space
@@
The next definition is based on [27, Lemma 4].
@@
and $r_{t_{1}}+r_{t_{2}}+\cdots+r_{t_{m(x,k)}}=r_{k}$, the linear map $\varsigma_{k}(x)\in\mathscr{M}_{k}|_{x}=End(\mathscr{D}_{x}^{(r_{k})})$ acts on each $s$th component
@@
[…]
@@
Since it is not hard to see that the boundary decomposition property is preserved under addition, multiplication, and involution, we only need to show the assertion when $\varsigma=\pi(g)$ for $g\in C(X)$ or $\varsigma=\pi(\xi)$ for $\xi\in\mathcal{E}_{Y}$.
@@
and the restriction $\varsigma_{k}(x)|_{\mathscr{V}_{x}^{(i)}}:\mathscr{V}_{x}^{(i)}\to\mathscr{V}_{x}^{(i)}$ is a multiplication by a number which we write as $\varsigma_{k}^{(i)}(x)$ as observed in (14) with $m=0$. On $s$th component
$\mathscr{D}^{(R_{x,k,s-1},r_{t_{s}})}_{x}=\big{(}\mathscr{V}_{\alpha^{R_{x,k,s-1}}(x)}^{(0)}\oplus\mathscr{V}_{\alpha^{R_{x,k,s-1}}(x)}^{(1)}\oplus\cdots\oplus\mathscr{V}_{\alpha^{R_{x,k,s-1}}(x)}^{(r_{t_{s}}-1)}\big{)}\otimes\mathscr{V}_{x}^{(R_{x,k,s-1})},$
@@
$\mathrm{diag}(g(\alpha^{R_{x,k,s-1}}(x)),g(\alpha^{R_{x,k,s-1}+1}(x)),\ldots,g(\alpha^{R_{x,k,s-1}+r_{t_{s}}-1}(x))).$
Whereas $\varsigma_{t_{s}}\in\Gamma(\mathscr{M}_{t_{s}})$ acts at $\alpha^{R_{x,k,s-1}}(x)\in Y_{t_{s}}$ as
@@
Now let $\varsigma=\varsigma_{1}\oplus\cdots\oplus\varsigma_{K}=\pi_{1}(\xi)\oplus\cdots\oplus\pi_{K}(\xi)$ for some $\xi\in\mathcal{E}_{Y}$. Then $\pi_{k}(\xi)(x)$ acts on the $s$th component of $\mathscr{D}_{x}^{(r_{k})}$ as follows. If we have $(a_{0},a_{1},\dots,a_{r_{k}-1})\in\mathscr{D}_{x}^{(r_{k})}$, then its $s$th component is equal to
@@
=\ = $\pi_{k}(\xi)(x)|_{\mathscr{D}_{x}^{(R_{x,k,s-1},r_{t_{s}})}}(a_{R_{x,k,s-1}},\dots,a_{R_{x,k,s}-1})$
=\ = $(0,\xi(\alpha^{R_{x,k,s-1}}(x))\otimes a_{R_{x,k,s-1}},\dots)$
=\ = $\pi_{t_{s}}(\xi)(\alpha^{R_{x,k,s-1}}(x))(a_{R_{x,k,s-1}},\dots,a_{R_{x,k,s}-1})$
=\ = $\varsigma_{t_{s}}(x)(\alpha^{R_{x,k,s-1}}(x))(a_{R_{x,k,s-1}},\dots,a_{R_{x,k,s}-1}),$
@@
Recall from (14) that we have a map
@@
for all $0\leq i\leq r_{k}-1,$ and vanishes on $\alpha^{-m}(Y)\cup\alpha^{-m+1}(Y)\cup\cdots\cup\alpha^{-1}(Y)$.
@@
where $\varsigma_{k}^{(i)}(x)=0$ for $i\geq r_{k}-m$. The equation (24) implies that for $i$ with $R_{x,k,s-1}\leq i<R_{x,k,s}$,
@@
for $i=0,1,2,\ldots,r_{k}-1$. $G$ is well-defined if $x\in\alpha^{i}(\overline{Y_{k}})\cap\alpha^{j}(\overline{Y_{l}})$ implies
@@
By (7), it is enough to show that if $x\in\alpha^{i}(\overline{Y_{k}})\cap\alpha^{j}(\overline{Y_{l}})\cap\alpha^{p}(Y_{q})$, then (25) holds. Note that
@@
whenever $0\leq i<r_{k}$ and $0\leq p<r_{q}$. Since otherwise, $0<p-i<r_{q}$, but we would have $\overline{Y_{k}}\cap\alpha^{p-i}(Y_{q})\neq\emptyset$, contradicting with the fact that $Y\cap\alpha^{p-i}(Y_{q})=\emptyset$.
@@
Let $\varsigma=(\varsigma_{1},\ldots,\varsigma_{K})\in\Gamma(\mathscr{M}_{1})\oplus\cdots\oplus\Gamma(\mathscr{M}_{K})$. Then $\varsigma\in\pi(C(X))$ if and only if $\varsigma$ has the boundary decomposition property, and $\varsigma_{k}(x)\in\Gamma(\mathscr{M}_{k}|_{x})$ has a diagonal matrix representation for all $x\in\overline{Y_{k}}$ and $k=1,\dots,K$.
@@
Since $\Gamma(\mathscr{M}_{1})\oplus\cdots\oplus\Gamma(\mathscr{M}_{K})$ is generated (as a $C^{*}$-algebra) by elements that have $m$th (lower) subdiagonal matrix representation, we obtain the following theorem.
@@
8.0.1. RSH structure of $\mathcal{O}(\mathcal{E}_{Y})$

### v2

@@ 5. $(e)$
Let $\varsigma=(\varsigma_{1},\ldots,\varsigma_{K})\in\Gamma(\mathscr{M}_{1})\oplus\cdots\oplus\Gamma(\mathscr{M}_{K})$. For an integer $10\leq m\leq r_{K}-$ 1, we say that $\varsigma_{k}\in\Gamma(\mathscr{M}_{k})$ has an $m$th subdiagonal matrix representation if at every $x\in\overline{Y_{k}}$, the matrix $(H_{\mathbf{U}}^{(k)})^{-1}(\varsigma_{k}(x))$ is $m$th subdiagonal for some (hence for all) $\mathbf{U}\in\mathcal{U}^{(r_{K})}$ with $x\in\mathbf{U}$. Here $(H_{\mathbf{U}}^{(k)})^{-1}(\varsigma_{k}(x))$ is regarded as the zero matrix if $m\geq r_{k}$.
Assume that each $\varsigma_{k}\in\Gamma(\mathscr{M}_{k})$ has an $m$th lower subdiagonal matrix representation. By (14), for $x\in\overline{Y_{k}}$, the map
@@
From the definition of $\pi_{k}$, we know that if $g\in C(X)$, then $\pi_{k}(g)(x)$ is represented as a diagonal matrix for all $x\in\overline{Y_{k}}$. Also if $\xi\in\mathcal{E}_{Y}^{\otimes m}\subset\mathcal{O}(\mathcal{E}_{Y})$ for some $11\leq m\leq r_{k}-$ 1, then $\pi_{k}(\xi)(x)$ is represented as an $m$th lower subdiagonal matrix since it maps $\mathscr{V}_{x}^{(i)}$ into $\mathscr{V}_{x}^{(i+m)}$ by (22). Also if $\varsigma=(\varsigma_{1},\ldots,\varsigma_{K})=\pi(\xi)$ for $\xi\in\mathcal{E}_{Y}^{\otimes m}$), then we have $\varsigma_{k}^{(i)}(x)=\psi(\xi)(\alpha^{i}(x))$ by (15) and (23).
@@
We assume $r_{k}>1$. Since the $C^{*}$-algebra $\Gamma(\mathscr{M}_{k})$ is generated by the sections $\varsigma\in\Gamma(\mathscr{M}_{k})$ which is an $m$th lower subdiagonal matrix in its matrix representation for $0\leq m<r_{k}$, it is enough to show that $\varsigma_{k}\in\pi_{k}(\mathcal{O}(\mathcal{E}_{Y}))$ for any such section $\varsigma_{k}$. Note that if $0\leq i\neq j<r_{k}$, then $\alpha^{i}(Y_{k})\cap\alpha^{j}(Y_{k})=\emptyset$ by definition of $r_{k}$.
Let $\varsigma_{k}\in\Gamma(\mathscr{M}_{k})$ have $m$th lower subdiagonal matrix representation. From the discussion given before this lemma, for $i=0,\dots,r_{k}-m-1$, there exists $\varsigma_{k}^{(i)}(x)\in\mathscr{V}_{\alpha^{i}(x)}^{(m)}$ such that the map
is given by $v\mapsto\varsigma_{k}^{(i)}(x)\otimes v$ for $v\in\mathscr{V}^{(i)}_{x}$.
@@
$\xi|_{\alpha^{i}(Y_{k})}(\alpha^{i}(x)):=\varsigma_{k}^{(i)}(x)$
for $x\in Y_{k}$ and $i=0,\dots,r_{k}-m-1$. This $\xi$ is well defined since the sets $\alpha^{i}(Y_{k})$ are mutually disjoint. It is easy to see that
@@
and then by Tietze extension theorem for vector bundles [5, Lemma 1.4.1], $\xi$ extends to a section $\widetilde{\xi}\in\Gamma(\mathscr{V}^{(m)})$ over the whole space $X$ such that for all $10\leq i\leq r_{k}-m-$ 1,
@@
$=$ $\ (0,\ldots,0,\varsigma_{k}^{(0)}(x)\otimes a_{0},\varsigma_{k}^{(1)}(x)\otimes a_{1},\ldots,\varsigma_{k}^{(r_{k}-m-1)}(x)\otimes a_{r_{k}-m-1})$
$=$ $\ \varsigma_{k}(x)(a_{0},a_{1},\ldots,a_{r_{k}-1}),$
hence $\pi_{k}(\psi^{-1}(\widetilde{\xi}))=\varsigma_{k}$ follows. ∎
@@ 8. A recursive subhomogeneous decomposition of $\mathcal{O}(\mathcal{E}_{Y})$
In this section we explicitly describe the image of $\mathcal{O}(\mathcal{E}_{Y})$ under the injective -homomorphism $\pi:\mathcal{O}(\mathcal{E}_{Y})\to\Gamma(\mathscr{M}_{1})\oplus\dots\oplus\Gamma(\mathscr{M}_{K})$. In particular, we construct a recursive subhomogeneous decomposition of $\pi(\mathcal{O}(\mathcal{E}_{Y}))\cong\mathcal{O}(\mathcal{E}_{Y})$. As in the case of $\mathrm{C}^{*}$-algebras of minimal homeomorphisms, we investigate the behaviour of boundary points. However, the presence of a twist (that is, a non-trivial line bundle) leads to further difficulties since we need to consider the local trivialization of the underlying endomorphism bundles when we are tracking the travel of boundary points through the Rokhlin towers.
@@
Note that the vector space
@@
The next definition is based on [30, Lemma 4].
@@
and $r_{t_{1}}+r_{t_{2}}+\cdots+r_{t_{m(x,k)}}=r_{k}$, the linear map $\varsigma_{k}(x)\in\mathscr{M}_{k}|_{x}=End(\mathscr{D}_{x}^{(r_{k})})$ acts on the $s$th component
@@
with respect to the local trivialization induced from $\{(U,h_{U})\mid U\in\mathcal{U}^{(1)}=\mathcal{U}\}$ of $\mathscr{V}$.
@@
Since it is not hard to see that the boundary decomposition property is preserved under addition, multiplication, involution, and taking limits, we only need to show the assertion when $\varsigma=\pi(g)$ for $g\in C(X)$ or $\varsigma=\pi(\xi)$ for $\xi\in\mathcal{E}_{Y}$.
@@
and the restriction $\varsigma_{k}(x)|_{\mathscr{V}_{x}^{(i)}}:\mathscr{V}_{x}^{(i)}\to\mathscr{V}_{x}^{(i)}$ is a multiplication by a number which we write as $\varsigma_{k}^{(i)}(x)$ as observed in (23) with $m=0$. On the $s$th component
$\mathscr{D}^{(R_{x,k,s-1},r_{t_{s}})}_{x}=\big{(}\mathscr{V}_{\alpha^{R_{x,k,s-1}}(x)}^{(0)}\oplus\mathscr{V}_{\alpha^{R_{x,k,s-1}}(x)}^{(1)}\oplus\cdots\oplus\mathscr{V}_{\alpha^{R_{x,k,s-1}}(x)}^{(r_{t_{s}}-1)}\big{)}\otimes\mathscr{V}_{x}^{(R_{x,k,s-1})}$
@@
$\mathrm{diag}(g(\alpha^{R_{x,k,s-1}}(x)),g(\alpha^{R_{x,k,s-1}+1}(x)),\ldots,g(\alpha^{R_{x,k,s-1}+r_{t_{s}}-1}(x))),$
whereas $\varsigma_{t_{s}}\in\Gamma(\mathscr{M}_{t_{s}})$ acts on $\alpha^{R_{x,k,s-1}}(x)\in Y_{t_{s}}$ as
@@
Now let $\varsigma=\varsigma_{1}\oplus\cdots\oplus\varsigma_{K}=\pi_{1}(\xi)\oplus\cdots\oplus\pi_{K}(\xi)$ for some $\xi\in\mathcal{E}_{Y}$. Then $\pi_{k}(\xi)(x)$ acts on the $s$th component of $\mathscr{D}_{x}^{(r_{k})}$ as follows. If $(a_{0},a_{1},\dots,a_{r_{k}-1})\in\mathscr{D}_{x}^{(r_{k})}$, then its $s$th component is equal to
@@
$=\pi_{k}(\xi)(x)|_{\mathscr{D}_{x}^{(R_{x,k,s-1},r_{t_{s}})}}(a_{R_{x,k,s-1}},\dots,a_{R_{x,k,s}-1})$
$=(0,\xi(\alpha^{R_{x,k,s-1}}(x))\otimes a_{R_{x,k,s-1}},\dots,\xi(\alpha^{R_{x,k,s}-2}(x))\otimes a_{R_{x,k,s}-2})$
$=\pi_{t_{s}}(\xi)(\alpha^{R_{x,k,s-1}}(x))(a_{R_{x,k,s-1}},\dots,a_{R_{x,k,s}-1})$
$=\varsigma_{t_{s}}(x)(\alpha^{R_{x,k,s-1}}(x))(a_{R_{x,k,s-1}},\dots,a_{R_{x,k,s}-1}),$
@@
Recall from (23) that if $\varsigma_{k}\in\Gamma(\mathscr{M}_{k})$ has an $m^{th}$ lower subdiagonal matrix representation, then we have a map
@@
for all $0\leq i\leq r_{k}-1,$ and vanishes on $\alpha^{-m}(Y)\cup\alpha^{-m+1}(Y)\cup\cdots\cup\alpha^{-1}(Y)$ provided $1\leq m<r_{k}$.
@@
where $\varsigma_{k}^{(i)}(x)=0$ for $i\geq r_{k}-m$. Equation (25) implies that for $i$ with $R_{x,k,s-1}\leq i<R_{x,k,s}$,
@@
for $i=0,1,2,\ldots,r_{k}-1$. The section $G$ is well defined if $x\in\alpha^{i}(\overline{Y_{k}})\cap\alpha^{j}(\overline{Y_{l}})$ implies
@@
By (7), it is enough to show that if $x\in\alpha^{i}(\overline{Y_{k}})\cap\alpha^{j}(\overline{Y_{l}})\cap\alpha^{p}(Y_{q})$, then (26) holds. If $0<p-i<r_{q}$, then we would have $\overline{Y_{k}}\cap\alpha^{p-i}(Y_{q})\neq\emptyset$, contradicting the fact that $Y\cap\alpha^{p-i}(Y_{q})=\emptyset$. Thus
@@
whenever $0\leq i<r_{k}$ and $0\leq p<r_{q}$.
@@
Let $\varsigma=(\varsigma_{1},\ldots,\varsigma_{K})\in\Gamma(\mathscr{M}_{1})\oplus\cdots\oplus\Gamma(\mathscr{M}_{K})$. Then $\varsigma\in\pi(C(X))$ if and only if $\varsigma$ has the boundary decomposition property, and $\varsigma_{k}(x)\in\Gamma(\mathscr{M}_{k}|_{x})$ has a diagonal matrix representation for all $x\in\overline{Y_{k}}$, $k=1,\dots,K$.
@@
Since $\Gamma(\mathscr{M}_{1})\oplus\cdots\oplus\Gamma(\mathscr{M}_{K})$ is generated (as a $\mathrm{C}^{*}$-algebra) by elements that have $m$th (lower) subdiagonal matrix representation, we obtain the following theorem.
@@
8.0.1. An RSH structure for $\mathcal{O}(\mathcal{E}_{Y})$